# UL-UNAS: Ultra-Lightweight U-Nets for Real-Time Speech Enhancement via Network Architecture Search

### v1

@@ 3.4 Loss function
Table 1: Performance of model prototypes equipped with various basic blocks.
@@
IDs Blocks Params (k) MACS (M/s) SISNR PESQ ESTOI (×$\times 100$) DNSMOS-P.835 DNSMOS-P.808
@@
Table 2: Ablation study results of the APReLU activation. The values in the PReLU and Affine columns represent the shapes of their corresponding learnable parameters, where “-” means term absence.
IDs PReLU Affine Params (k) MACS (M/s) SISNR PESQ ESTOI (×$\times 100$) DNSMOS-P.835 DNSMOS-P.808
@@
1 $1$ 37.23 23.72 10.80 1.99 74.08 2.59 2.94 3.81 3.46
2 $C$ 37.50 23.72 10.91 2.01 74.47 2.58 2.94 3.78 3.45
3 $C\times F$ 49.86 23.72 10.97 2.02 74.41 2.60 2.95 3.81 3.48
4 $1$ $1$ 37.27 23.72 10.75 1.99 74.04 2.60 2.94 3.81 3.47
5 $C$ $C$ 38.08 23.72 10.67 2.00 73.96 2.60 2.95 3.81 3.47
6 $C$ $C\times F$ 62.79 23.72 10.97 2.04 74.50 2.63 2.97 3.84 3.50
7 $C\times F$ $C\times F$ 75.14 23.72 11.01 2.05 74.57 2.61 2.95 3.81 3.50
4 Experiments
4.1 Datasets
[…]
4.2 Implementation details
4.2.1 Parameter configurations for model prototype
The BM module adopts the same setup in GTCRN, which maps the 192 high-frequency bands to 64 ERB bands, while keeping the 65 low-frequency bands unaltered, leading to a 129-dimensional compressed feature map. We use 5 identical blocks from the basic blocks described in Sec. 3.2.1 or the extended blocks described in Sec. 3.2.4 to construct the encoder. All blocks have 16 channels and a kernel size of (3,3). The first two blocks have a stride of 2 in the frequency dimension to downsample the feature map, while the remaining three blocks have a stride of 1. The decoder is a mirror version of the encoder, where all the standard or depthwise convolutional layers are replaced with transposed ones with the same parameter configurations.
4.2.2 Training configurations
The STFT is computed using a Hanning window with a length of 32 ms, a hop length of 16 ms, and an FFT size of 512 points. The log-power spectrogram is fed into the model, which outputs an estimated IRM to obtain the enhanced magnitude. The enhanced speech is then obtained by combining the enhanced magnitude with the noisy phase. The models are trained using the Adam optimizer (Kingma and Ba, 2015) with two different schedulers. For the DNS3 dataset, a linear-warm-up cosine-annealing scheduler is used. The learning rate increases linearly from $10^{-6}$ to $10^{-3}$ over the first 25,000 steps and then decays cosine-wise until 250,000 steps, at which point the training procedure ends. A batch size of 32 is used for the full DNS3 dataset, while a batch size of 8 is used for the DNS3 subset. For the VCTK-DEMAND dataset, a performance-based learning rate scheduler is applied. The learning rate starts at $10^{-3}$ and is halved if the validation loss fails to decrease for 5 consecutive epochs. For this dataset, a batch size of 4 is used.
4.2.3 Neural architecture search configurations
The controller cell has an embedding size of 100, and an LSTM hidden size of 200. It is trained using the Adam optimizer with an initial learning rate of 0.001. The multi-processing technique is employed to sample and evaluate multiple architectures concurrently. We sample 40 architectures per episode and train them on 8 RTX 3090 GPUs, with each GPU running 5 models. Each model is trained on the DNS3 subset for 20 epochs with 5,000 noisy-clean pairs randomly selected at each epoch to serve as the training data. Validation is conducted only during the final three epochs, and the average perceptual evaluation of speech quality (PESQ) (Recommendation, 2001) score is used as the $Q(m)$ in Eq. (11). The baseline score $Q_{0}$ is set to 1, and the target MACS $M_{T}$ is set to 30M. The two weight constants $\omega_{+}$ and $\omega_{-}$ are set to -0.15 and 0, respectively. If the reward does not increase for 5 consecutive epochs, the learning rate is halved.
4.3 Models for comparison and evaluation metrics
We compare our model with the latest SOTA ultra-lightweight SE models, including GTCRN (Rong et al., 2024) and LiSenNet (Yan et al., 2024). When comparing on the VCTK-DEMAND dataset, some models from this benchmark are also used for comparison, including DeepFilterNet (Schröter et al., 2022), FullSubNet (Hao et al., 2021), FastFullSubNet (Hao and Li, 2022), and GaGNet (Li et al., 2022). We use the statistics reported in the original papers if available; otherwise, we retrain these models using the official implementations with our training configurations to ensure fairness.
We use three intrusive metrics for performance evaluation: the scale-invariant signal-to-noise ratio (SISNR) (Roux et al., 2019), the PESQ, and the extended short-time objective intelligibility (ESTOI) (Jensen and Taal, 2016). Additionally, two non-intrusive metrics, DNSMOS P.808 (Reddy et al., 2021b) and DNSMOS P.835 (Reddy et al., 2022), are employed to provide a more comprehensive comparison. DNSMOS P.835 includes three sub-metrics: OVRL for overall speech quality, SIG for signal quality, and BAK for background noise quality.
Table 3: Ablation study results of the cTFA module.
IDs FITA FDTA Params (k) MACS (M/s) SISNR PESQ ESTOI (×$\times 100$) DNSMOS-P.835 DNSMOS-P.808
@@ 4.3 Models for comparison and evaluation metrics
3 37.64 25.14 10.78 1.99 74.08 2.60 2.94 3.82 3.47
4 85.63 28.29 11.52 2.12 75.98 2.64 2.98 3.83 3.51
5 Results and analysis
5.1 Ablation study
5.1.1 Effect of basic blocks
To explore the potential of various convolutional blocks, we compare the performance of different model prototypes, each equipped with distinct convolutional blocks. The results are presented in Table 1. It can be observed that when the encoder uses only Conv blocks, it achieves fairly good performance, but at the cost of a significantly heavier computational load, as is shown in ID1. On the other hand, using efficient convolutional blocks provides a trade-off between computation and performance. Among the blocks with two convolutional layers (from ID2 to ID4), the DWS block stands out for its significant reduction in computational cost, albeit with an inevitable decline in performance. In contrast, the Ghost block achieves a marginal performance advantage but at the expense of the highest computational cost. Meanwhile, the Rep block, which employs a reparameterization technique, fails to deliver meaningful performance improvements despite its added complexity. Given these observations, we conclude that the DWS block offers a more efficient balance between performance and computational cost. In the case of blocks with three convolutional layers (ID5 and ID6), the MB block demonstrates a clear performance advantage with less MACS. Based on these results, the Conv block, DWS block, and MB block are selected as representatives for subsequent experiments.
5.1.2 Effect of affine PReLU
In this section, we investigate the application of APReLU to the DWS block to demonstrate its effectiveness. Specifically, We design the following experiments: (1) analyzing the impact of the slope-controlling parameter shape in PReLU, and (2) examining the effect of the shapes of the two affine parameters. The results are presented in Table 2. Note that the computational overhead of affine transformations is disregarded, as it is minimal and negligible.
It can be observed that extending the shape of the slope-controlling parameter leads to a slight improvement in model performance, as shown from ID1 to ID3. We attribute this to more fine-grained activations across different dimensions. Similarly, the performance boost from the affine transformation is also closely tied to the parameter shapes. When the affine parameters are frequency-independent, the gains are negligible, as shown in ID4 and ID5. However, when the parameters are frequency-dependent, more substantial performance gains are observed, as shown in ID6 and ID7. Comparing ID6 and ID7, the improvements in intrusive metrics are minimal, and the DNSMOS scores even degrade. This suggests that using both a frequency-dependent slope-controlling parameter and affine parameters appears somewhat redundant. Overall, the APReLU activation provides substantial performance gains, with a 0.05 improvement in PESQ and a 0.04 increase in OVRL and DNSMOS-P.808.
@@ 5.1.2 Effect of affine PReLU
5.1.3 Effect of causal time-frequency attention
In this section, we explore the application of cTFA to the DWS block to demonstrate its effectiveness. Specifically, we design the following experiments: (1) employing only the FITA module, (2) employing only the FDTA module, and (3) employing both the FITA and FDTA modules. The results are presented in Table 3, which show that the FITA module significantly enhances the model’s performance, achieving a 0.1 increase in PESQ, as observed in ID1 and ID2. In contrast, the exclusive application of the FDTA module provides little improvement, as shown in ID3. However, by integrating both attention branches, the model achieves the best performance, with a total PESQ improvement of 0.13, as demonstrated in ID4. This suggests that the frequency-dependent characteristics of the FDTA module complement the FITA module. Notably, while both the FITA and FDTA modules introduce additional parameters, they incur negligible computational overhead.
Table 4: Optimal network configuration discovered by NAS.
@@ 5.1.3 Effect of causal time-frequency attention
Table 5: Performance of models equipped with diverse combinations of blocks.
IDs Blocks Params (k) MACS (M/s) SISNR PESQ ESTOI (×$\times 100$) DNSMOS-P.835 DNSMOS-P.808
@@
1 XConv x 5 112.21 61.81 11.87 2.20 76.94 2.68 3.01 3.85 3.50
2 XDWS x 5 111.18 26.70 11.66 2.15 76.33 2.65 2.99 3.83 3.52
3 XMB x 5 114.18 33.64 11.80 2.19 77.00 2.66 3.00 3.84 3.51
4 NAS-discovered 169.00 33.62 12.02 2.24 77.31 2.68 3.01 3.86 3.55
Table 6: Performance comparison on the DNS3 test set. LA represents look ahead. The values in BOLD indicate the best results in each metric.
Models LA (ms) Params (k) MACS (M/s) SISNR PESQ ESTOI (×$\times 100$) DNSMOS-P.835 DNSMOS-P.808
@@
GTCRN 0 48$\star$ 34$\star$ 11.44 2.10 75.71 2.63 2.97 3.82 3.50
LiSenNet 32 37 65 12.12 2.23 77.02 2.70 3.04 3.85 3.55
UL-UNAS (proposed) 0 169 34 12.25 2.28 77.92 2.70 3.04 3.88 3.56
$\star$ We adopt the updated statistics from the official GitHub repository of GTCRN.
Table 7: Performance comparison on the VCTK-DEMAND test set. The values in BOLD indicate the best results in each metric.
@@
5.1.4 Effect of neural architecture search
In this section, we first explore the application of both APReLU and cTFA to the three basic blocks, forming extended blocks that serve as candidate blocks for NAS. We present the trend of reward changes with respect to the number of sampled models in Fig. 8. It can be observed that the reward of the top-1 model improves steadily as the number of sampled models increases. After sampling more than 1200 models, the rewards of the top-1, top-5, and top-25 models reach close proximity, indicating that the NAS process is nearing convergence. The entire process lasts for approximately 2 days. The optimal network configuration discovered by NAS is presented in Table 4. The performance results of the discovered model along with models equipped with the three extended blocks are shown in Table 5. By comparing ID1, ID2, and ID5 in Table 1 with ID1, ID2, and ID3 in Table 5, we observe that the application of both APReLU and cTFA leads to substantial improvements across all metrics, highlighting the broad effectiveness of both modules. Moreover, the NAS technique further enhances performance by leveraging different block combinations. Interestingly, NAS tends to favor models with a larger number of parameters, even though we impose no restrictions on the parameter count. This somewhat suggests the importance of having more parameters in improving model performance. The architecture discovered by NAS achieves superior performance compared to the overall best manually designed prototype network (ID3) while maintaining slightly fewer MACS. Notably, due to a reduced total number of layers, the NAS-discovered architecture also results in lower inference latency.
5.2 Comparison with baselines
@@
For a more comprehensive comparison with other models, we also train UL-UNAS on the VCTK-DEMAND dataset, with the results presented in Table 7. Notably, we provide two sets of performance results for UL-UNAS: one from the model trained with the standard loss function and the other from the model trained with an additional PESQ loss [^1] The PESQ loss is implemented by torch-pesq packet: https://github.com/audiolabs/torch-pesq.. This approach ensures alignment with the results reported in the original LiSenNet paper and facilitates a fair comparison. It is evident that UL-UNAS outperforms LiSenNet in both cases. Furthermore, UL-UNAS demonstrates the best performance among all the models while maintaining lower computational costs and a relatively compact model size. Audio examples are available at https://github.com/Xiaobin-Rong/ul-unas.
6 Conclusions
In this paper, we present UL-UNAS, an ultra-lightweight SE model that requires only 34M MACS. By systematically exploring efficient convolutional blocks and introducing innovative components including the APReLU and cTFA modules, we construct a more powerful U-Net family that excels in computational efficiency and enhancement quality. Furthermore, the employment of NAS enables the discovery of an optimized architecture tailored to real-time constraints. Through these contributions, UL-UNAS not only surpasses SOTA ultra-lightweight models but also demonstrates competitive results against more resource-intensive baseline models, making it a robust solution for real-time SE applications.
7 Acknowledgement
This work was supported by the National Nature Science Foundation of China (Grant No. 12274221) and the AI & AI for Science Project of Nanjing University.

### v2

@@ IV-C Baselines and evaluation metrics
TABLE I: Performance of model prototypes equipped with various basic blocks.
@@
IDs Blocks Params (k) MACS (M/s) SISNR PESQ ESTOI ($\times 100$) DNSMOS-P.835 DNSMOS-P.808
@@
TABLE II: Ablation study results of the APReLU activation and comparison with other activation functions. The values in parentheses under the “Act.” and “Affine” columns specify the dimensionality of the corresponding parameters.
IDs Act. Affine Params (k) MACS (M/s) SISNR PESQ ESTOI ($\times 100$) DNSMOS-P.835 DNSMOS-P.808
@@
1 PReLU (1) 37.23 23.72 10.80 1.99 74.08 2.59 2.94 3.81 3.46
2 PReLU (C) 37.50 23.72 10.91 2.01 74.47 2.58 2.94 3.78 3.45
3 PReLU (CxF) 49.86 23.72 10.97 2.02 74.41 2.60 2.95 3.81 3.48
4 PReLU (1) ✓ (1) 37.27 23.72 10.75 1.99 74.04 2.60 2.94 3.81 3.47
5 PReLU (C) ✓ (C) 38.08 23.72 10.67 2.00 73.96 2.60 2.95 3.81 3.47
6 PReLU (C) ✓ (CxF) 62.79 23.72 10.97 2.04 74.50 2.63 2.97 3.84 3.50
7 PReLU (CxF) ✓ (CxF) 75.14 23.72 11.01 2.05 74.57 2.61 2.95 3.81 3.50
8 GELU 37.22 23.72 10.87 1.99 74.02 2.60 2.94 3.82 3.47
9 GELU ✓ (CxF) 62.50 23.72 10.94 2.03 74.35 2.62 2.96 3.83 3.50
10 Swish 37.50 22.92 10.76 1.99 74.05 2.60 2.94 3.82 3.47
11 Swish ✓ (CxF) 62.79 22.92 10.97 2.03 74.35 2.62 2.96 3.82 3.50
TABLE III: Ablation study results of the cTFA module.
IDs TA FA Params (k) MACS (M/s) SISNR PESQ ESTOI ($\times 100$) DNSMOS-P.835 DNSMOS-P.808
@@
3 39.99 26.43 10.85 2.00 74.12 2.61 2.95 3.82 3.48
4 87.98 29.59 11.66 2.15 76.20 2.65 2.99 3.85 3.53
@@
TABLE IV: Optimal network configuration discovered by NAS.
@@
TABLE V: Performance of models equipped with diverse combinations of blocks.
IDs Blocks Params (k) MACS (M/s) SISNR PESQ ESTOI ($\times 100$) DNSMOS-P.835 DNSMOS-P.808
@@
1 XConv $\times$ 5 114.56 63.11 11.96 2.22 77.08 2.69 3.02 3.86 3.53
2 XDWS $\times$ 5 113.53 28.00 11.69 2.16 76.47 2.66 2.99 3.84 3.53
3 XMB $\times$ 5 116.53 34.94 11.85 2.20 77.01 2.67 3.00 3.84 3.53
4 NAS-discovered 171.33 34.91 12.07 2.25 77.69 2.69 3.01 3.86 3.55
TABLE VI: Performance comparison on the DNS3 test set. LA represents look ahead. All models are retrained with the same training configuration as UL-UNAS.
Models LA (ms) Params (k) MACS (M/s) SISNR PESQ ESTOI ($\times 100$) DNSMOS-P.835 DNSMOS-P.808
@@
CRN-small 0 195 35 11.34 2.07 75.22 2.62 2.95 3.84 3.49
DPCRN-small 0 15 38 10.67 2.01 74.29 2.59 2.94 3.84 3.49
DeepFilterNet 0 1780 358 12.20 2.13 75.92 2.73 3.07 3.86 3.43
GTCRN 0 48 34 11.44 2.10 75.71 2.63 2.97 3.82 3.50
LiSenNet 32 37 56 12.12 2.23 77.02 2.70 3.04 3.85 3.55
UL-UNAS (proposed) 0 171 35 12.30 2.30 78.24 2.71 3.05 3.88 3.56
TABLE VII: Performance comparison on the VCTK-DEMAND test set. The values in BOLD indicate the best results in each metric.
@@
V Results and analysis
V-A Ablation study
V-A1 Effect of base blocks
To explore the potential of various convolutional blocks, we compare the performance of different model prototypes, each equipped with distinct convolutional blocks. The results are presented in Table I. It can be observed that when the encoder uses only Conv blocks, it achieves fairly good performance, but at the cost of a significantly heavier computational load, as is shown in ID1. On the other hand, using efficient convolutional blocks provides a trade-off between computation and performance. Among the blocks with two convolutional layers (from ID2 to ID4), the DWS block stands out for its significant reduction in computational cost, albeit with an inevitable decline in performance. In contrast, the Ghost block achieves a marginal performance advantage but at the expense of the highest computational cost. Meanwhile, the Rep block, which employs a reparameterization technique, fails to deliver meaningful performance improvements despite its additional operational complexity. Given these observations, we conclude that the DWS block offers a more efficient balance between performance and computational cost. In the case of blocks with three convolutional layers (ID5 and ID6), the MB block demonstrates a clear performance advantage with fewer MACS. Based on these results, the Conv block, DWS block, and MB block are selected as representatives for subsequent experiments.
V-A2 Effect of affine PReLU
In this section, we investigate the application of APReLU to the DWS block to demonstrate its effectiveness. Specifically, we design the following experiments: (1) analyzing the impact of the slope-controlling parameter shape in PReLU, and (2) examining the effect of the shapes of the two affine parameters. The results are presented in Table II. Note that the computational overhead of affine transformations is disregarded, as it is minimal and negligible.
It can be observed that extending the shape of the slope-controlling parameter yields slight performance improvements from ID1 to ID3, likely due to more fine-grained activations across different dimensions. Similarly, the performance boost from the affine transformation is also closely tied to the parameter shapes. When the affine parameters are frequency-independent, the gains are negligible, as shown in ID4 and ID5. However, when the parameters are frequency-dependent, more noticeable performance gains are observed, as shown in ID6 and ID7, highlighting the importance of frequency-dependent affine recalibration. Comparing ID6 and ID7, the improvements in intrusive metrics are minimal, and the DNSMOS scores even degrade. This suggests that using both a frequency-dependent slope-controlling parameter and affine parameters appears somewhat redundant. Overall, the APReLU activation provides consistent and meaningful performance gains, with a 0.05 improvement in PESQ and a 0.04 increase in OVRL and DNSMOS-P.808.
To enable a comprehensive comparison, we also incorporate existing advanced activation functions, including GELU [14] and Swish [36]. As shown by ID8 and ID10 versus ID6, APReLU consistently achieves superior performance. Furthermore, applying our affine transformation to GELU (ID9) and Swish (ID11) also yields measurable gains (PESQ: +0.04, DNSMOS OVRL: +0.02). This result confirms that the frequency-dependent affine recalibration is a general-purpose design principle, consistent with its effectiveness on PReLU.
V-A3 Effect of causal time-frequency attention
In this section, we explore the application of cTFA to the DWS block to demonstrate its effectiveness. Specifically, we design the following experiments: (1) employing only the TA module, (2) employing only the FA module, and (3) employing both the TA and FA modules. The results are presented in Table III, which shows that the TA module significantly enhances the model’s performance, achieving a 0.1 increase in PESQ, as observed in ID1 and ID2. In contrast, the exclusive application of the FA module provides little improvement, as shown in ID3. However, by integrating both attention branches, the model achieves the best performance, with a total PESQ improvement of 0.16, as demonstrated in ID4. This suggests that the frequency-dependent characteristics of the FA module complement the TA module. Notably, while both the TA and FA modules introduce additional parameters, the resulting computational overhead remains modest and acceptable.
V-A4 Effect of neural architecture search
@@
V-B Comparison with baselines
[…]
For a more comprehensive comparison with other models, we also train UL-UNAS on the VCTK-DEMAND dataset, with the results presented in Table VII. Notably, we provide two sets of performance results for UL-UNAS: one from the model trained with the standard loss function and the other from the model trained with an additional PESQ loss [^2] The PESQ loss is implemented by torch-pesq packet: https://github.com/audiolabs/torch-pesq.. This approach ensures alignment with the results reported in the original LiSenNet paper and facilitates a fair comparison. It is evident that UL-UNAS outperforms LiSenNet in both cases. Furthermore, UL-UNAS demonstrates the best performance among all the models while maintaining lower computational costs and a relatively compact model size.
VI Conclusions
In this paper, we present UL-UNAS, an ultra-lightweight SE model that requires only 35M MACS. By systematically exploring efficient convolutional blocks and introducing innovative components, including the APReLU and cTFA modules, we construct a more powerful U-Net family that excels in computational efficiency and enhancement quality. Furthermore, the employment of NAS enables the discovery of an optimized architecture tailored to real-time constraints in a short search time, presenting a practical and scalable approach for automating lightweight model design. Through these contributions, UL-UNAS not only surpasses SOTA ultra-lightweight models but also demonstrates competitive results against more resource-intensive baseline models, making it a robust solution for real-time SE applications.